OpenStudy (anonymous):

if 56 C flows past a point in 10.0 A of current, then how much time elapsed? 2.0 s , 5.6 s , 11.2 s , or 20.0 s ? :/

OpenStudy (michele_laino):

the requested itime interval is: \Delta t = Q/I= 56/10=...seconds

OpenStudy (anonymous):

oh! 5.6! so choice B? :O

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!! :)

OpenStudy (michele_laino):

:)